To verify Cayley-Hamilton theorem for

1 2 3

M = 2 -1 4

3 1 1


1
Expert's answer
2021-03-02T06:00:17-0500

"\mid A- \lambda I=0\\ equation 1
- \lambda^3+\lambda^2+18\lambda+30=0\\

- \lambda^3+\lambda^2+18\lambda+30I=0\\

Calculate \, A^2=\begin{bmatrix}
 14 & 3&14\\
 12 & 9&6\\
8&6&14\\

\end{bmatrix}\\
Calculate \, A^3=
\begin{bmatrix}
 62& 39&68\\
 48 & 21&78\\
62&24&62\\
\end{bmatrix}
\\Put \, all \, values \, in \, equation \, 1\\\\=



 
 


 
 
\begin{bmatrix}
 -62& -39&-68\\
 -48 & -21&-78\\
 -62&-24&-62
\end{bmatrix}+
\begin{bmatrix}
 14& 3&14\\
 12 & 9&6\\
 8&6&14\\

\end{bmatrix}+18

\begin{bmatrix}
 1& 2&3 \\
 2& -1&4\\
 3&1&1\\
\end{bmatrix}+30
\begin{bmatrix}
 1& 0&0 \\
 0& 1&0\\
 0&0&1\\
\end{bmatrix}\\
=


\begin{bmatrix}
 0&0&0\\
 0&0&0\\
0&0&0
\end{bmatrix}\\
L.H.s=R.H.s\\
So\, this\, matrix\, meets \,Caley \, Hamilton\, Theorem"